a) Juan is riding his bike eastward at a speed of 15 km/h and heads the siren of an ambulance (700 Hz) that moves west at 60 km/h. Which frequency will the child heard?

b) A sound wave propagates through a 10 cm closed tube. The air temperature is -8 Celsius degrees. Calculate the second and third harmonics.


1
Expert's answer
2021-09-29T09:55:22-0400

a) The frequency the boy will hear is


"f=\frac{330+15·1000/3600}{330-60·1000/3600}·700=747\text{ Hz}."

b) The speed of sound is


"v=\sqrt{\gamma\frac{RT}{M}}=326\text{ m/s}."


The second and third harmonics are


"f_2=\frac{n_2v}{2L}=3260\text{ Hz},\\\space\\
f_3=\frac{n_3v}{2L}=4890\text{ Hz}."